# On the Design and Analysis of LLM-Based Algorithms

### v1

@@ 4 Conclusion and discussion
We have introduced a formal framework for studying the design and analysis of LLM-based algorithms. After identifying the computational-graph representation, the principle of task decomposition, and other abstractions, we find it feasible to provide formal analysis for the accuracy and efficiency of generic LLM-based algorithms. As a case study, we provide in-depth analysis for the basic pattern of parallel decomposition and four concrete examples of it, which is further validated by extensive numerical experiments.
Moving forward, we find it promising future research to further expand the proposed framework and thereby advance the field of LLM-based algorithms, from a theoretical or practical perspective. For example:
@@
  • While our proposed framework is, in principle, applicable to generic LLM-based algorithms, our detailed investigation of concrete examples has been limited to the basic pattern of parallel decomposition. It would be interesting to apply our framework to more diverse patterns of LLM-based algorithms, e.g. deeper sequential decomposition or divide-and-conquer algorithms with recursive decomposition, which can potentially bring new insights for algorithm design.
  • In our empirical study, we have mainly focused on the hyperparameter $m$ that indicates the granularity of parallel task decomposition. Future work might place more emphasis on other configurations of LLM-based algorithms, such as the choices of LLM models, prompting methods (e.g. chain-of-thoughts or step-by-step reasoning), LLM decoding methods (e.g. greedy vs. stochastic [CDH+24]), among others.
@@
We believe that there are plenty of opportunities for future research on LLM-based algorithms, and would like to invite the community to contribute to this exciting and rapidly developing field, using the current work as a starting point.
@@ Acknowledgements
The authors would like to thank the AgentScope team, especially Xuchen Pan, Dawei Gao and Weirui Kuang, for their assistance with using AgentScope and setting up LLM inference service.

### v2

@@ 7 Conclusion and discussion
We have introduced a formal framework for studying the design and analysis of LLM-based algorithms. After identifying the computational-graph representation, the principle of task decomposition, and other abstractions, we find it feasible to provide formal analysis for the accuracy and efficiency of generic LLM-based algorithms. In multiple case studies, we provide in-depth analysis with numerical validation for various patterns of LLM-based algorithms, including parallel, hierarchical and recursive decomposition.
Moving forward, we find it promising future research to further expand or apply the proposed framework and thereby advance the field of LLM-based algorithms, from a theoretical or practical perspective. For example:
@@
  • In our empirical study, we have mainly focused on the hyperparameter $m$ that indicates the granularity of parallel task decomposition, or the option of prompting certain LLM calls to either answer directly or think step by step. Future work might place more emphasis on other configurations of LLM-based algorithms, such as the choices of LLM models (e.g. choosing a strong-but-expensive model for specific LLM calls and a weak-but-cheap one for the others within the same algorithm), LLM decoding methods (e.g. greedy vs. stochastic [WWS+23, CDH+24, BJE+24]), etc., and better understand how they impact the accuracy and efficiency of LLM-based algorithms.
@@
  • From a more practical perspective, future work might adopt the proposed framework, or some components and methodology within it, to assist the design, analysis, improvement and application of new LLM-based algorithms, or for fair comparison between different algorithms, with both accuracy and efficiency taken into account.
Given the plenty of opportunities for future research on LLM-based algorithms, we would like to invite the community to contribute to this exciting and rapidly developing field, using the current work as a starting point.
@@ Acknowledgements
The authors would like to thank Yuchang Sun for helpful discussion, as well as the AgentScope team, especially Xuchen Pan, Dawei Gao and Weirui Kuang, for their assistance with using AgentScope and setting up LLM inference service for the experiments.